# envname-P envname#1

### v1

@@ 4. Design and Implementation
Cargo Sherlock is a Python-based tool for quantifying trust in the Rust supply chain ecosystem for a given Rust crate. It provides a logs mode, which collects and displays information about the crate, and a trust mode (see output in Figure˜1), which is used to compute the trust cost and severity label of the crate, with components as detailed in the following subsections.
Figure 2. High level overview of Cargo Sherlock.:
@@ 4.1. Data collection
Cargo Sherlock first gathers crate metadata from multiple sources. From crates.io, it collects information about the author(s), download counts, dependencies, and GitHub repository stars and forks. The audit reports are sourced from Cargo-Vet (mozilla2024cargovet, ) and include organizations such as Google, Mozilla, and Bytecode Alliance. It also gathers the outputs from static analysis tools while additional details on unsoundness, and vulnerabilities are gathered using Miri (miri, ) and RustSec (rustsec, ).
@@ 4.2. Relational model
The aggregated data is organized into a relational model as shown in Figure˜3. The primary entities include: Crates, Authors, Audits, Dependencies, and Tool Results. Each Crate includes basic metadata such as its name, version, downloads, and related git data (e.g., stars and forks). Crates can have multiple associated tool results, audits, authors, and dependencies, following a one-to-many relationship pattern. Dependencies connect crates to other crates, while Audits link organizations and trust criteria to crates.
Figure 3. The Cargo Sherlock relational model (ER diagram).:
@@ 4.3. Selection of assumptions
Our candidate assumptions are motivated by prior work on software supply chain trust. Hamer et al. (hamer2025trusting, ) identified key trust factors among Rust developers, while Boughton et al. (boughton2024decomposing, ) decomposed these into distinct dimensions. Table˜2 shows some of the assumptions we support and their justifications in prior research.
Cargo Sherlock supports three kinds of assumptions: positive, parameterized, and negative. The costs assigned to each candidate assumption are configurable by the user. In our implementation, we provide default costs which we arrived at through a process of trial and error.
@@
25–100 If the crate has many downloads, it is safe. (hamer2025trusting, ) (Tbl. 7).
5 If the current version has a passed audit, it is safe. (hamer2025trusting, ) (Tbl. 7).
20 If a past version has a passed audit, it is safe. (hamer2025trusting, ) (Tbl. 7).
20–100 If the repository has many stars/forks, it is safe. (hamer2025trusting, ) (Tbl. 7).
10 If the crate has no side effects and all dependencies are safe, it is safe. (safe-haskell, ; cargo-scan-paper, )
5 If the author of the crate is trusted, it is safe. (hamer2025trusting, ) (Tbl. 7).
@@
5 If the crate appears on RustSec with a critical severity label, it is not safe. (boughton2024decomposing, )
@@
90 If the crate appears on RustSec and the provided crate version is patched, it is not safe. (boughton2024decomposing, )
30 If the crate is flagged by Miri, it is not safe. (boughton2024decomposing, )
60–100 If the crate has many side effects, it is not safe. (safe-haskell, ; cargo-scan-paper, )
10 If the crate has an unsafe dependency, it is not safe. (boughton2024decomposing, )
Table 2. Selected trust and distrust model assumptions with justification and costs.
@@
Positive Assumptions: These are candidate assumptions a user might use to reason that a crate is safe. By default, Cargo Sherlock includes the following:
@@
  • Trusted Author: A crate is safe if it is authored by a well-established developer in the Rust community (hamer2025trusting, ). Cargo Sherlock includes a default list of trusted developers, which users can customize.
@@
  • Audit Passed: A crate is safe if it or previous versions of it have passed an audit (hamer2025trusting, ).
@@
  • Side Effects and Safe Dependencies: A crate is safe if it has no side effects, i.e., no observable change to program state or the external environment beyond returning a value, and all of its dependencies are safe. (safe-haskell, ; cargo-scan-paper, ).
@@
  • Crate is Safe: Cargo Sherlock supports assuming that a crate is safe with no justification as the base case. This assumption has the highest cost associated with it.
Parameterized Assumptions: These are the candidate assumptions for which the trust cost is calculated as a function of certain metadata parameters; we use a smoothing function so that the trust cost changes gradually with changes to the parameter. By default, Cargo Sherlock includes the following:
  • Number of Downloads: A crate with more downloads is safe (hamer2025trusting, ).
  • GitHub Stars and Forks: A crate with more stars and forks on its GitHub is safe (hamer2025trusting, ).
@@
Negative Assumptions: These are candidate assumptions a user might use to reason that a crate is unsafe. By default, Cargo Sherlock includes the following:
  • RustSec: A crate is unsafe if it appears in the RustSec advisory database (rustsecAdvisoriesx203A, ).
@@
  • Tool Results: Cargo Sherlock supports the integration of static analysis tools. In our current setup, Cargo Sherlock supports the assumptions that crates flagged by Miri (miri, ) are unsafe (boughton2024decomposing, ) and crates with many side effects identified by cargo-scan (cargo-scan, ) are unsafe (cargo-scan-paper, ; safe-haskell, ). The latter is a parameterized assumption.
@@
Cargo Sherlock also applies a consistency checker to ensure some checks on the assumption costs; currently, it checks that no assumption has a cost higher than the base cost of 100.
4.4. Computing the trust score
@@
We use the Z3 SMT solver (z3, ) to compute minimum trust and distrust costs using Algorithm˜1 and Algorithm˜2 (default). In our implementation, we replace the explicit binary search over costs with Z3’s optimizing SAT interface to obtain the minimum cost directly. The trust cost ranges from 0 to 100, where a lower trust cost represents a crate that is easier to trust. Similarly, the distrust cost ranges from 0 to 100, where a lower distrust cost represents a crate that is easier to distrust. The final label is derived by combining both costs, using the function illustrated in Figure˜4.
0202040406060808010010002020404060608080100100SafeLow SeverityMedium SeverityHigh SeverityCriticalSafeUnsafeDistrust CostTrust Cost
@@
Figure 4. Cargo Sherlock combines trust and distrust scores to produce a severity label.:
@@ 5. Evaluation
  • RQ4: How does the performance of Algorithm˜1 and Algorithm˜2 scale with the number of dependencies?
All evaluations were conducted on an M2-Pro MacBook Pro with 16 GB of RAM. Our primary experiments focused on two sets of crates: the 100 most downloaded crates from crates.io and a random sample of 1000 crates. Additionally, we included crates flagged in the RustSec database to measure the tool’s performance on known vulnerable crates. [^2] All experiments were run between March 2025 and November 2025.
@@
For all experiments, we adopt an idealized rational-developer model in the sense of Dimoulas et al. (dimoulas2025rational, ). The developer interacts with our tool as follows: when they intend to add a dependency to their existing Rust project, they execute Cargo Sherlock on the selected crate. This generates a comprehensive report accompanied by a severity label; we assume that the developer looks at the report and decides whether to go forward with adding the dependency based on the severity label. For each experimental section, we interpret our findings through this user model and discuss the resulting implications for prospective users of Cargo Sherlock.
@@ 5.1. RQ1: Does Cargo Sherlock detect synthetically generated typosquatting attacks?
Typosquatting is an attack technique that exploits typographical errors to deceive users into downloading a malicious package, such as publishing serde_yml in place of serde_yaml. We consider typosquatting attacks because it is common type of supply chain attack, and align closely with Cargo Sherlock’s intended use case. To address RQ1, we constructed synthetic typosquatted versions of the top 100 most frequently downloaded Rust crates from crates.io. We selected popular crates as a test sample, because their popularity makes them a lucrative target for an attacker. We generated typosquatted variants by modifying the name and author of each crate, while keeping the functionality, dependencies and side effects unchanged. To determine how many downloads would be obtained by a lesser-known crate by default, we put dummy crates up on crates.io. We found within 3 months, the crates received an average of 171 downloads. We use this number for the download count in our synthetic experiment. After creating these typosquatted variants, we evaluated them using Cargo Sherlock and compared their severity labels with corresponding original crates. Figure˜5(a) shows the distribution of obtained severity labels. Most of the original crates are classified as safe (73%) and no crate is classified higher than low_severity. On the other hand, among typosquatted crates, most are classified as critical. However, a few typosquatted crates also get the safe label because they have no side effects and they all have safe dependencies. The severity label increases for the typosquatted crate when compared to the original crate in all cases.
RQ1 Findings: All typosquatted versions of popular Rust crates incur substantially higher trust costs and higher severity labels compared to their original counterparts when evaluated by Cargo Sherlock. Implication: A user interacting with the tool would identify the typosquatted crate variant when adding a dependency to their code.
(a) Heat map showing the severity label distribution for real and synthetic typosquatted crates as assigned by Cargo Sherlock
Crate Trust Cost Distrust Cost Severity Level
Pair 2: fast_log (original) vs faster_log (typosquatted)
@@
Pair 1: serde_yaml (original) vs serde_yml (AI-maintained fork)
@@
(b) Summary table showing the trust, distrust cost and the severity label assigned to actual and typosquatted version crate
Figure 5. Overview of severity results: (a) Section˜5.1 synthetic experiment (b) Section˜5.2 real world example:
@@ 5.2. RQ2: Does Cargo Sherlock detect real-world supply chain risks?
We address this question as a case study on Rust crates that have been involved in known incidents on crates.io. Of such crates that exist to our knowledge, we were able to obtain the source code of two of them for our analysis. [^3] Unfortunately, we were unable to obtain the source code for the typosquatted versions of rustdecimal (rustsecRUSTSEC20220042Rustdecimal, ) and asyncprintln (fasterlog-asyncprintln, ) to include them in this analysis.
We first examine fast_log and faster_log; faster_log was identified as a malicious typosquat of fast_log (fasterlog-asyncprintln, ): beyond reproducing the original crate’s logging functionality, it injects code that scans log files for sensitive material (including cryptocurrency private keys) and exfiltrates any matches to a remote endpoint. Cargo Sherlock assigns a MEDIUM_SEVERITY label (trust cost $37$, distrust cost $73$) to fast_log, basing the trust on the author (they have a history of publishing crates) and the distrust on the presence of side effects (it performs file I/O). In contrast, Cargo Sherlock assigns faster_log a CRITICAL label, with a substantially higher trust cost ($75$) and elevated distrust cost ($80$). The change in authorship (the author had no other crates) and side effects leads to a different trust and distrust cost.
Our second case involves serde_yaml, authored by David Tolnay, a well-established and trusted developer in the Rust community. After serde_yaml was marked as unmaintained, a fork, serde_yml (tolnay2024impression, ), was published on crates.io and had accumulated over 1,057,000 downloads by March 19, 2025. The fork incorporates a substantial amount of AI-generated code, including low-quality (“AI slop”) changes that were later called out by Tolnay (tolnay2024impression, ). As shown in Table 5(b), the original crate receives a SAFE label, while the fork is classified as LOW_SEVERITY, primarily due to the change in authorship. The distrust costs are similar across the two crates since no malicious code or side effects were introduced.
RQ2 Findings: Cargo Sherlock successfully assigns higher severity labels to the malicious typosquat faster_log and AI-maintained variant serde_yml, relative to their original crates. Implication: Users interacting with Cargo Sherlock may perceive these crates as risky.
@@ 5.3. RQ3: How does Cargo Sherlock output fare on known vulnerable crates?
To address RQ3, we compare Cargo Sherlock’s severity labels with those from RustSec (rustsec, ), a well-known database of security advisories for Rust crates. RustSec classifies crates by severity levels: critical, high, medium, low, and informational. We analyzed 592 unique crate-version pairs listed in RustSec [^4] This is as of March 25, 2025. For this experiment, all negative assumptions related to RustSec were removed to determine what labels would have been assigned to vulnerable crates prior to them being known to be vulnerable.
Figure˜6(a) illustrates the distribution of severity labels for crates listed on RustSec. The x-axis shows the severity labels assigned by RustSec and each bar is split according to the labels assigned by Cargo Sherlock. While many crates receive a higher severity label, most crates are classified as SAFE or LOW_SEVERITY. These crates are typically popular. Cargo Sherlock classified 61 out of 592 crates as SAFE-many of these crates have an INFO label on RustSec, often indicating that the crate is unmaintained, but they still exhibit high engagement metrics (such as high download count). Figure˜6(b) confirms that this behavior is driven by the popularity of RustSec-listed crates: approximately 75% of RustSec crates belong to the top 10% of crates.io by downloads, and only around 15% of RustSec crates are found in the bottom 80%. In contrast, HIGH_SEVERITY and CRITICAL labels correspond to crates with numerous side effects, failed MIRI tests, low download counts, relatively unknown authors, and a lack of audits.
(a) Distribution of Cargo Sherlock severity labels on RustSec data with RustSec labels.
@@
Figure 6. Comparison of severity label distributions: (a) RustSec crates and (b) crates in RustSec by downloads.:
@@
5.4. RQ4: How does the performance of Algorithm˜1 and Algorithm˜2 scale with the number of dependencies?
To address RQ4, we sampled 1,000 random crates from crates.io. We excluded 47 crates from our evaluation because all of their versions had been yanked, leaving 953 crates. To isolate algorithmic running time, we excluded data-collection time from all measurements. The cumulative distribution function (CDF) plot in Figure˜7(a) shows the runtime for both algorithms, using a timeout of 600 seconds. Algorithm˜1 successfully evaluated 298 crates and timed out on 500 crates. It crashes on 155 crates, out of which 150 crashes are because of max recursion limit reached on the system. The remaining 5 crashes are shared between both algorithms and are due to parsing issues. Algorithm˜2 successfully evaluated 900 crates and only timed out on 48 crates.
(a) CDF of running time for Algorithm˜1 vs. Algorithm˜2.
(b) Scatter plot of time taken vs. number of dependencies.
Figure 7. Results of scalability experiments on 1000 random crates.:
The scatter plot (Figure˜7(b)) shows that Algorithm˜1 rapidly becomes infeasible as the number of dependencies increases, with most evaluations timing out beyond approximately 10 dependencies. Conversely, Algorithm˜2 algorithm scales better, successfully evaluating crates with up to 150 dependencies. A regression fit suggests that the scaling of the Horn clause algorithm is exponential. This is expected, as unfold in Algorithm˜2 generates an exponential DNF formula in the worst case.
RQ4 Findings: Algorithm˜1 becomes infeasible after 10 dependencies, while Algorithm˜2 scales to at least 150 dependencies. Implication: Users can run Cargo Sherlock on crates with a large number of dependencies and receive results within 10 minutes.

### v2

@@ 4. Design and Implementation
Cargo Sherlock is a Python-based tool for quantifying trust in the Rust supply chain ecosystem, given as input a Rust crate. [^3] https://github.com/davispl/cargo-sherlock It provides a logs mode, which collects and displays information about the crate, and a trust mode (see output in Figure 1), which is used to compute the trust cost and severity label of the crate, with components as detailed in the following subsections.
Figure 2. High level overview of Cargo Sherlock.
High level overview of the Cargo Sherlock tool. The tool gathers information from multiple sources, builds a relational model, and uses Z3 solver to compute the trust cost and severity label of a given Rust crate.
@@ 4.1. Data collection
Cargo Sherlock first gathers crate metadata from multiple sources. From crates.io, it collects information about the author(s), download counts, dependencies, and GitHub repository stars and forks. From Cargo Vet (Mozilla, 2024), it sources audit reports from organizations including Google, Mozilla, and Bytecode Alliance. Cargo Sherlock also gathers the outputs from static analysis tools (Miri (Rust team, 2024) and Cargo Scan (PL and PLSysSec, 2026; Zoghbi et al., 2026)) and vulnerability information from RustSec (Rust Secure Code Working Group, 2025).
@@ 4.2. Relational model
The aggregated data is organized into a relational model as shown in Figure 3. The primary entities include: Crates, Authors, Audits, Dependencies, and Tool Results. Each Crate includes basic metadata such as its name, version, downloads, and related git data (e.g., stars and forks). Crates can have multiple associated tool results, audits, authors, and dependencies, following a one-to-many pattern. Dependencies connect crates to other crates, while Audits link organizations and trust criteria to crates.
Refer to caption Entity-Relationship (ER) diagram of the Cargo Sherlock relational model, showing entities such as Crate, Author, Audit, Dependency, and Tool Results, along with their attributes and relationships.
Figure 3. Cargo Sherlock entity-relationship (ER) model.
@@ 4.3. Selection of assumptions
Our candidate assumptions are motivated by prior work on software supply chain trust. Hamer et al. (Hamer et al., 2025) identified key trust factors among Rust developers, while Boughton et al. (Boughton et al., 2024) decomposed these into distinct dimensions. Table 2 shows some of the assumptions we support and their justifications in prior research.
Cargo Sherlock supports three kinds of assumptions: positive, negative, and parameterized. The costs assigned to each candidate assumption are configurable by the user. In our implementation, we provide default costs, which we arrived at through a process of trial and error.
@@
25–100 If the crate has many downloads, it is safe. (Hamer et al., 2025) (Tbl. 7)
5 If the crate version has a passed audit, it is safe. (Hamer et al., 2025) (Tbl. 7)
20 If a past version has a passed audit, it is safe. (Hamer et al., 2025) (Tbl. 7)
20–100 If the repository has many stars/forks, it is safe. (Hamer et al., 2025) (Tbl. 7)
10 If the crate has no side effects and all dependencies are safe, it is safe. (Terei et al., 2012; Zoghbi et al., 2026)
5 If the author of the crate is trusted, it is safe. (Hamer et al., 2025) (Tbl. 7)
@@
5 If the crate appears on RustSec with a critical severity label, it is not safe. (Boughton et al., 2024)
@@
90 If the crate appears on RustSec and the provided crate version is patched, it is not safe. (Boughton et al., 2024)
30 If the crate is flagged by Miri, it is not safe. (Boughton et al., 2024)
60–100 If the crate has many side effects, it is not safe. (Terei et al., 2012; Zoghbi et al., 2026)
10 If the crate has an unsafe dependency, it is not safe. (Boughton et al., 2024)
Table 2. Selected assumptions used by Cargo Sherlock (including justifications in prior work).
Side-by-side tables summarizing trust and distrust assumptions with costs and justifications.
Positive assumptions:
@@
These are candidate assumptions a user might use to reason that a crate is safe. By default, Cargo Sherlock includes the following:
@@
  • Trusted Author: A crate is safe if it is authored by a well-established developer in the Rust community (Hamer et al., 2025). Cargo Sherlock includes a default list of trusted developers, which users can customize.
@@
  • Audit Passed: A crate is safe if it has passed an audit (Hamer et al., 2025). This assumption has a variant for crates with a previous version that have passed an audit.
@@
  • Side Effects and Safe Dependencies: A crate is safe if it has no side effects, i.e., no observable change to program state or the external environment beyond returning a value, and all of its dependencies are safe (Terei et al., 2012; Zoghbi et al., 2026). Side effects are identified using Cargo Scan (PL and PLSysSec, 2026).
@@
  • Crate is Safe: By default, a crate is safe (this assumption has the highest possible cost). Cargo Sherlock includes this assumption with no justification as a base case.
Negative assumptions:
@@
These are candidate assumptions a user might use to reason that a crate is unsafe. By default, Cargo Sherlock includes the following:
  • RustSec: A crate is unsafe if it appears in the RustSec advisory database (Rust Secure Code Working Group, 2025). This assumption has variants for crates with a previous version that appear in RustSec and for different RustSec labels.
@@
  • Tool Results: A crate is unsafe if it is flagged by a static analysis tool (Boughton et al., 2024). Currently, Cargo Sherlock supports the assumption that crates flagged by Miri (Rust team, 2024) are unsafe.
  • Unsafe Dependency: A crate is unsafe if it has an unsafe dependency. (Boughton et al., 2024)
  • Crate is Unsafe: By default, a crate is unsafe (this assumption has the highest possible cost). Cargo Sherlock includes this assumption with no justification as a base case.
Parameterized assumptions:
These are candidate assumptions for which the trust cost is calculated as a function of certain metadata parameters; we use a smoothing function so that the trust cost changes gradually with changes to the parameter. We also allow parameterized negative assumptions, which work for distrust cost in the same way. By default, Cargo Sherlock includes the following:
  • Number of Downloads: A crate with many downloads is safe (Hamer et al., 2025).
  • GitHub Stars and Forks: A crate with many stars and forks on its GitHub is safe (Hamer et al., 2025).
  • Side Effects: A crate with many side effects is unsafe (Terei et al., 2012; Zoghbi et al., 2026).
@@
Cargo Sherlock applies a consistency checker to ensure checks on all assumption costs; currently, it checks that no assumption has a cost higher than the base cost of 100.
4.4. Computing the trust cost
@@
We use the Z3 SMT solver (de Moura and Bjørner, 2008) to compute minimum trust and distrust costs using either Algorithm 1 or Algorithm 2 (default). In our implementation, we replace the explicit binary search over costs with Z3’s optimizing SAT interface to obtain the minimum cost directly. The trust cost ranges from 0 to 100, where a lower trust cost represents a crate that is easier to trust. Similarly, the distrust cost ranges from 0 to 100, where a lower distrust cost represents a crate that is easier to distrust. The final label is derived by combining both costs, using the function illustrated in Figure 4.
@@
Figure 4. Cargo Sherlock combines trust and distrust costs to produce a severity label.
A scatter plot showing the relationship between distrust cost and trust cost, divided into regions labeled Safe, Low Severity, Medium Severity, High Severity, and Critical. The plot illustrates how Cargo Sherlock combines these costs to produce a severity label for Rust crates.
@@ 5. Evaluation
  • RQ4: How does the performance of Algorithms 1 and 2 scale with the number of dependencies?
Experimental setup.
All evaluations were conducted on an M2-Pro MacBook Pro with 16 GB of RAM. Our experiments are run on various sets of crates as described below: the 100 most downloaded crates from crates.io, a random sample of 1000 crates, a sample of crates flagged in the RustSec database, and two case studies. All experiments were run between March and November 2025.
@@
For all experiments, we adopt an idealized “rational developer” model in the sense of Dimoulas et al. (Dimoulas and Felleisen, 2025). The rational developer interacts with our tool as follows: when they intend to add a dependency to their existing Rust project, they execute Cargo Sherlock on the selected crate. This generates a report accompanied by a severity label; we assume that the developer looks at the report and decides whether to go forward with adding the dependency based on the severity label. For each experimental section, we interpret our findings through this user model and discuss the resulting implications for prospective users of Cargo Sherlock.
@@ 5.1. RQ1: Does Cargo Sherlock detect synthetically generated typosquatting attacks?
Typosquatting is an attack technique that exploits typographical errors to deceive users into downloading a malicious package, such as downloading serde_yml in place of serde_yaml. We consider typosquatting because it is a common category of supply chain attack, and aligns closely with Cargo Sherlock’s intended use case. To address RQ1, we constructed synthetic typosquatted versions of the top 100 most frequently downloaded Rust crates from crates.io. We selected popular crates as a test sample because their popularity makes them a lucrative target for an attacker. We generated typosquatted variants by modifying the name and author of each crate, while keeping the functionality unchanged. To determine how many downloads would be obtained by a lesser-known crate by default, we measured that placeholder crates on crates.io received an average of 171 downloads within 3 months. We use this number for the download count in our synthetic experiment. After creating typosquatted variants, we evaluated them using Cargo Sherlock and compared their severity labels with corresponding original crates. Figure 5(a) shows the distribution of obtained severity labels. Most of the original crates are classified as safe (73%) and no crate is classified higher than low_severity. On the other hand, among typosquatted crates, most are classified as critical. However, 25 typosquatted crates also receive the safe label, likely because they have no side effects and they all have safe dependencies. The severity label increases for the typosquatted crate when compared to the original crate in all other cases.
RQ1 Findings: Most typosquatted versions of popular Rust crates incur substantially higher trust costs and higher severity labels compared to their original counterparts when evaluated by Cargo Sherlock. Implication: A user interacting with the tool would identify the typosquatted crate variant when adding a dependency to their code.
(a) Cargo Sherlock results on synthetically generated typosquatting attack crates (compared to their real versions).
Cargo Sherlock severity labels on synthetically generated typosquat crates (compared to their real versions).
Crate Trust Cost Distrust Cost Severity Label
Pair 1: fast_log (original) vs faster_log (typosquatted)
@@
Pair 2: serde_yaml (original) vs serde_yml (AI-maintained fork)
@@
(b) Cargo Sherlock results on real-world incidents.
A table summarizing trust cost, distrust cost, and severity level for two pairs of crates: fast_log vs faster_log, and serde_yaml vs serde_yml.
Figure 5. Sections 5.1 and 5.2 experimental results.
RQ1 and RQ2 severity results overview.
@@ 5.2. RQ2: Does Cargo Sherlock detect real-world supply chain risks?
We address this question as a case study on Rust crates that have been involved in known incidents on crates.io. We obtained the source code for two such crates that exist to our knowledge: faster_log and serde_yml. [^4] We are aware of known incidents for at least two other Rust crates: rustdecimal (Rust Secure Code Working Group, 2024b) and asyncprintln (Pearce, 2025). Unfortunately, we were unable to obtain the source code for the typosquatted versions of these crates to include them in the analysis.
We first examine faster_log, an identified malicious typosquat of fast_log (Pearce, 2025). Beyond reproducing the original crate’s logging functionality, it injects code that scans log files for sensitive material (including cryptocurrency private keys) and exfiltrates any matches to a remote endpoint. Cargo Sherlock assigns a MEDIUM_SEVERITY label (trust cost $37$, distrust cost $73$) to fast_log, basing the trust on the author (they have a history of publishing crates) and the distrust on the presence of side effects (it performs file I/O). In contrast, Cargo Sherlock assigns faster_log a CRITICAL label, with a substantially higher trust cost ($75$) and elevated distrust cost ($80$). The change in authorship (the author had no other crates) and side effects leads to a different trust and distrust cost.
Our second case is serde_yml, a fork of serde_yaml (the latter authored by David Tolnay, a well-established and trusted developer in the Rust community). After the original crate was marked unmaintained, the fork was published on crates.io and had accumulated over 1,057,000 downloads by March 19, 2025 (Tolnay, 2025). The fork incorporates a substantial amount of AI-generated code, including low-quality “AI slop” changes that were later criticized by Tolnay (Tolnay, 2025). As shown in Table 5(b), the original crate receives a SAFE label, while the fork is classified with the higher LOW_SEVERITY, primarily due to the change in authorship. The distrust costs are similar across the two crates since no identifiable malicious code or side effects were introduced.
RQ2 Findings: Cargo Sherlock assigns higher severity labels to the malicious typosquat faster_log and AI-maintained fork serde_yml, relative to their original crates. Implication: Users interacting with Cargo Sherlock may perceive these crates as risky.
@@ 5.3. RQ3: How does Cargo Sherlock output fare on known vulnerable crates?
To address RQ3, we compare Cargo Sherlock’s severity labels with those from RustSec (Rust Secure Code Working Group, 2025), a well-known database of security advisories for Rust crates. RustSec classifies crates by severity levels: critical, high, medium, low, and informational. We analyzed 592 unique crate-version pairs listed in RustSec. [^5] This number is as of March 25, 2025. For this experiment, all negative assumptions related to RustSec were removed to determine what labels would have been assigned to vulnerable crates prior to them being known to be vulnerable.
Figure 6(a) illustrates the distribution of severity labels for crates listed on RustSec. The x-axis shows the severity labels assigned by RustSec, and each bar is split according to the labels assigned by Cargo Sherlock. While some crates receive a higher severity label, most crates are classified as SAFE or LOW_SEVERITY. For example, these include popular crates (hence, more likely to be marked SAFE) that are listed in RustSec with an INFO label due to being marked as unmaintained. Figure 6(b) confirms that this behavior is generally driven by the popularity of RustSec-listed crates: approximately 75% of RustSec crates belong to the top 10% of crates.io by downloads, and only around 15% of RustSec crates are found in the bottom 80%.
(a) Distribution of Cargo Sherlock vs. RustSec severity labels.
A bar chart showing the distribution of Cargo Sherlock versus RustSec severity labels on the crates listed in RustSec.
@@
A bar chart showing the percentile distribution of crates in RustSec by downloads on crates.io.
Figure 6. Section 5.3 experimental results.
RQ3 results plots.
@@
5.4. RQ4: How does the performance of Algorithms 1 and 2 scale with the number of dependencies?
To address RQ4, we sampled 1,000 random crates from crates.io. We excluded 47 crates from our evaluation because all of their versions had been yanked, leaving 953 crates. To isolate algorithmic running time, we excluded data collection time from all measurements. The cumulative distribution function (CDF) plot in Figure 7(a) shows the runtime for both algorithms, using a timeout of 600 seconds. Algorithm 1 successfully evaluates 298 crates and timed out on 500 crates. It crashes on 155 crates, out of which 150 crashes are due to the system’s maximum recursion depth being exceeded. The remaining 5 crashes are shared by both algorithms and are due to parsing failures. Algorithm 2 successfully evaluated 900 crates and only timed out on 48 crates.
(a) CDF of running time for Algorithm 1 vs. Algorithm 2.
[…]
Figure 7. Section 5.4 scalability results.
RQ4 results plots.
The scatter plot (Figure 7(b)) shows that Algorithm 1 rapidly becomes infeasible as the number of dependencies increases, with most evaluations timing out beyond approximately 10 dependencies. Conversely, Algorithm 2 scales better, successfully evaluating crates with up to 150 dependencies. A regression fit suggests that the scaling of the Horn clause algorithm is exponential. This is expected, as unfold in Algorithm 2 generates an exponential DNF formula in the worst case.
RQ4 Findings: Algorithm 1 becomes infeasible after 10 dependencies, while Algorithm 2 scales to at least 150 dependencies. Implication: Users can run Cargo Sherlock on crates with a large number of dependencies and receive results within 10 minutes.